At 10 ˚C, the solubility of KNO3 is 30 g/100 g H2O. Would you expect the solubility of KNO3 to be higher or lower at 40 ˚C? Explain.

A sample containing 80 g of oxygen gas has a volume of 50 L. What is the volume after 4.00 g of oxygen gas is added to the balloon, if temperature and pressure do not change?
Answer #1

Solution:

Part A) Solubility of KNO3 increases with temperature. The increase in temperature causes an increase in kinetic energy. Thus, increase kinetic energy causes breaking of the solute by the solvent molecules more effectively. Thus, the solubility of KNO3 at 40 °C is more than solubility at 30 °C.

Part B) When temperature and pressure is kept constant, tgen volume is proportional to number of moles.

V1/V2 = n1/n2

n1 = number of moles = mass / molar mass

n1 = 80 g / 32 g mol-1 = 2.5 mol

n2 = 80 g + 4 g / 32 g.mol-1 = 84 g./ 32 g mol-1

n2 = 2.625 mol

Thus,

V1/V2 = n1/n2

50 L / V2 = 2.5 mol/ 2.625 mol

V2 = 50 L x 2.625 mol./ 2.5 mol = 52.5 L

Hence, volume = 52.5 L